Rolling a regular die -
Number:
1 (number of times rolled: 6)
2 (number of times rolled: 4)
3 (number of times rolled: 1)
4 (number of times rolled: 5)
5 (number of times rolled: 2)
6 (number of times rolled: ?)
Based on the histogram above,
1. How many sixes are rolled? 3 - is this correct?

2. What is the experimental probability P(2)? I got 4/6 - is this correct?

3. What is the experimental probability P(4)? 5/6 - is this correct?
Thank you


Answer by greenestamps(13200)



It appears you are misunderstanding the given data....

A regular die was rolled multiple times. It came up 1 6 times, 2 4 times, 3 1 time, 4 5 times, 5 2 times, and 6 an unknown number of times.

None of the three questions can be answered with the given information. We either need to know the number of times a 6 came up, or we need to know the total number of times the die was rolled, so that we can determine the number of times a 6 came up.

1. If your answer of 3 is correct, then the total number of rolls was 6+4+1+5+2+3 = 21. Was that total number of rolls given, and you forgot to include that information in your post?

For 2 and 3, the experimental probability is the number of times a certain number came up, divided by the total number of rolls.

Assuming the total number of rolls was in fact 21...

2. P(2) = 4/21
3. P(4) = 5/21

----------------------------------------------------

The reader responded saying that the die was rolled 25 times.

So, since the number of rolls other than 6 is 18, the answer to question 1 is that a 6 was rolled 7 times.

And then the answers to questions 2 and 3 are 4/25 and 5/25 = 1/5.